Can you pleae help me with these two questions:
The table of ordered pairs shown was generated by a graphing calculator
a. What is the x intercept
b. Find the slope of the line.
x y
0 1
-2 0
-4 -1

Answer by venugopalramana(3286)

X intercept of a line is the distance (to be measured along the positive direction of the X axis) from origin to the point where the line meets the X axis .As the intersecting point is on the X axis ,it is characterised by the fact that its Y coordinate is zero.From the table we find that when Y =0 , X = -2 . Hence the X intercept is -2.
Slope of a line is given by the formula
Difference in Y coordinates of any 2 points on the line / Difference in X coordinates of the same 2 points on the line .
If we take the 2 points as A (0,1) and B (-2,0) from the given ordered pairs...
slope = (0-1)/(-2-0)= -1/(-2)=1/2